### v1

# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

Abstract
Missing data is a pervasive problem in tabular settings. Existing solutions range from simple averaging to complex generative adversarial networks. However, due to huge variance in performance across real-world domains and time-consuming hyperparameter tuning, no default imputation method exists. Building on TabPFN, a recent tabular foundation model for supervised learning, we propose TabImpute, a pre-trained transformer that delivers accurate and fast zero-shot imputations requiring no fitting or hyperparameter tuning at inference-time. To train and evaluate TabImpute, we introduce (i) an entry-wise featurization for tabular settings, which enables a $100\times$ speedup over the previous TabPFN imputation method, (ii) a synthetic training data generation pipeline incorporating realistic missingness patterns, which boosts test-time performance, and (iii) MissBench, a comprehensive benchmark for evaluation of imputation methods with $42$ OpenML datasets and $13$ missingness patterns. MissBench spans domains such as medicine, finance, and engineering, showcasing TabImpute’s robust performance compared to $11$ established imputation methods.
1 Introduction
Missing data is ubiquitous across tabular datasets, affecting statisticians, economists, health officials, and businesses. For example, healthcare datasets may lack some recorded blood pressure measurements, or datasets merged from multiple sources may only share partial features. Regardless of the source, missing data must be imputed to numerical values before employing statistical or machine learning models.
Refer to caption Refer to caption
(a) Imputation Accuracy ($\uparrow$) (b) Runtime ($\downarrow$)
Figure 1: Evaluation on real-world OpenML data: MissBench. We compare TabImpute and TabImpute+ (ensembled method) with $11$ other popular methods on MissBench. In panel (a), we plot the imputation accuracy (defined as 1 - normalized RMSE), which is calculated for each method, normalized within a dataset, and averaged across datasets and $13$ missingness patterns. Error bars indicate 95% confidence intervals. In panel (b), we compare the runtime per table entry. Any method not labeled (GPU) is tested on a CPU because that method is not GPU-compatible. TabPFN on CPU is significantly slower, so we do not include it. See Sec. 3 for our exact computing specifications and Sec. 4 for accuracy score methodology.
Rubin (1976) proposed analyzing missingness based on its relationship with the data-generating process to determine whether missingness biases downstream estimation. Rubin demonstrated that when missingness operates independently of the underlying data, the observed data distribution provides an unbiased foundation for estimation. This framework categorizes missingness into three classes: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR) (Van Buuren, 2012; Sportisse et al., 2020a; b). MCAR defines scenarios where missingness is independent of observed and missing values; MAR denotes settings where missingness depends on observed variables; and MNAR—most general of them all—denotes settings where missingness can arbitrarily depend on observed and unobserved variables.
Imputation methods range from simple approaches (e.g., constant default values and averages) to more sophisticated techniques like nearest-neighbor-based methods (Batista & Monard, 2003) and random forest regression, notably the MissForest algorithm (Stekhoven & Bühlmann, 2011). However, each method in the literature is typically tailored for specific settings, with performance varying significantly across datasets, domains, and missingness patterns (Van Buuren, 2012; Jarrett et al., 2022; Agarwal et al., 2023; Ibrahim et al., 2005). Another class of methods come from matrix completion including USVT (Chatterjee, 2015), SoftImpute (Hastie et al., 2015), and nearest neighbor-based methods (Agarwal et al., 2023; Dwivedi et al., 2022; Chin et al., 2025). While USVT and SoftImpute were typically designed for MCAR settings, nearest neighbor-based methods are designed for complex MNAR settings (Chin et al., 2025).
Building on recent advances in tabular representation learning (Hollmann et al., 2023; Ye et al., 2025), we propose a pre-trained transformer specifically designed for the tabular missing-data problem that produces accurate and fast zero-shot imputations. TabPFN is a pre-trained transformer model for supervised learning that performs well across a variety of domains without any fine-tuning (Hollmann et al., 2025). The team behind TabPFN created an imputation method in their tabpfn-extensions Python package by using the TabPFN model in iterative column-wise imputation (available on GitHub [^1] https://github.com/PriorLabs/tabpfn-extensions/blob/main/src/tabpfn_extensions/unsupervised/unsupervised.py).
Starting point of this work is a comprehensive evaluation of the pre-existing methods, and we do so by building a new benchmark MissBench–consisting several real-world OpenML datasets and realistic missingness patterns induced on them. Next, we leverage the TabPFN architecture to introduce a new method TabImpute, which achieves state-of-the-art-performance on this benchmark. More specifically, our contributions can be summarized as follows:
We improve on this approach by introducing a new entry-wise featurization (EWF), allowing parallel prediction of each missing value using TabPFN’s model, denoted EWF-TabPFN. To further improve on TabPFN’s model for missing data imputation, we train a new underlying model, TabImpute, to better fit this class of tasks, outperforming fine-tuning TabPFN (shown in Sec. 4). Finally, we adaptively ensemble EWF-TabPFN’s and TabImpute’s predictions to achieve state-of-the-art results in imputation accuracy on our comprehensive benchmark. We denote the final ensembled method as TabImpute+.
MCAR
Seq-MNAR
Block-MNAR
Polarization-MNAR
Panel-MNAR
Figure 2: Selection of synthetic missingness patterns implemented in MissBench. Blue entries indicate observed values, and gray entries are unobserved.
The main contributions of this work can be summarized as follows:
  • We introduce a novel, comprehensive test bench, denoted MissBench, using $42$ real-world OpenML datasets and $13$ missingness patterns (see Sec. 4 for details). Our benchmark builds on previous work, namely HyperImpute (Jarrett et al., 2022) comprising of $13$ UC Irvine (UCI) datasets (Kelly et al., 2024), a subset of OpenML, with $3$ missingness patterns, and GAIN (Yoon et al., 2018) comprising of $6$ UCI datasets with $1$ missingness pattern.
  • Next we introduce a state-of-the-art method in following steps:
    • First, we propose a new entry-wise missing data featurization (see Sec. 3.1 for details). Notably, this new featurization, when directly used with the pre-trained TabPFN model of Hollmann et al. (2023), provides a significant improvement in speed and accuracy over TabPFN. We call this method EWF-TabPFN.
    • Next, we develop a synthetic data generation pipeline to create training datasets with missing values covering a wide range of missingness patterns (see Sec. 3.2, Sec. A.2, and Tab. 6 for details). We pre-train a TabPFN architecture on this pipeline with our new featurization—we call this method TabImpute. This model provides an improvement over EWF-TabPFN across several missingness patterns.
    • Finally, we introduce an adaptive ensembling imputation method, denoted TabImpute+, to combine the power of TabImpute and EWF-TabPFN (see Sec. 3.4 for details). This allows us to leverage the underlying world-knowledge of EWF-TabPFN and domain-knowledge for missing data from TabImpute to yield accurate imputations across heterogeneous datasets.
Our code and implementation details for all our contributions above can be accessed on GitHub. [^2] https://github.com/jacobf18/tabular
Figure 3: Overview of our contributions. The first row demonstrates TabPFN’s imputation method, which performs iterative column-by-column imputation. We build on this by introducing an entry-wise featurization, as shown in the second row. We create a new synthetic data-generator for missingness data to train our model, TabImpute, shown in green (Sec. 3.2 and Sec. 3.3, respectively). Lastly, we ensemble TabImpute with TabPFN’s model using our features to create TabImpute+ (Sec. 3.4). We adaptively evaluate all the imputers on the comprehensive and rich set of OpenML datasets with many missingness patterns applied (Sec. 4).
1.1 Previous Work
We build primarily on missing data imputation and tabular representation learning (TRL). While missing data imputation is well-studied with established theory dating back to the 1970s (Rubin, 1976), TRL using tabular foundation models is relatively new (Müller et al., 2022; Zhang et al., 2025; Hollmann et al., 2023; 2025). Below, we describe relevant work that we directly compare against or build upon.
Imputation methods.
Given the widespread nature of missing data, numerous imputation techniques have been proposed. These include averages, linear models over columns (Efron, 1994), random forest models (Hindy et al., 2024), ensemble methods (Jarrett et al., 2022), nearest neighbor-based methods (Chin et al., 2025), and even using generative adversarial networks (Yoon et al., 2018). For fully numerical data, matrix completion methods like SoftImpute (Hastie et al., 2015) have also been employed.
HyperImpute (Jarrett et al., 2022) combines the power of multiple classical imputation methods through iterative imputation (IM). IM loops over each column, using other columns to predict missing values until convergence. HyperImpute optimizes the imputer at each iteration over candidate methods, including MICE (Royston & White, 2011), SoftImpute (Hastie et al., 2015), column mean, MissForest (Stekhoven & Bühlmann, 2011), and Optimal Transport (Muzellec et al., 2020). While achieving state-of-the-art performance across diverse missingness patterns and supporting categorical variables, HyperImpute is designed for supervised learning settings and cannot handle matrices with entirely missing columns (common in causal inference and panel data (Agarwal et al., 2023)). Additionally, optimization requires significantly more time than base methods.
Tabular representation learning.
Tabular representation learning focuses on building models that can generalize across diverse tabular domains. The pioneering work in this area is TabPFN, a tabular foundation model for supervised learning. TabPFN was proposed in Hollmann et al. (2023) and subsequently improved in Hollmann et al. (2025). Since TabPFN’s introduction, numerous variants have emerged to address scalability and performance limitations, as well as further work aimed at clarifying its internal representations (Zhang et al., 2025; Ye et al., 2025). Recent advances include TabICL (Qu et al., 2025), a scalable foundation model that extends supervised learning capabilities to datasets with up to 500K samples through a novel two-stage architecture with column-then-row attention mechanisms. Other notable models include MITRA (Zhang & Robinson, 2025), a tabular foundation model pre-trained on purely synthetic data from a mix of random classifiers/regressors, and Toto (Cohen et al., 2025), which is optimized for time series forecasting on observability metrics. Additionally, CausalFM (Ma et al., 2025) enables Bayesian causal inference through structural causal model priors similar to TabPFN, while DO-PFN (Robertson et al., 2025) extends the PFN framework to estimate counterfactual distributions from observational data.
2 Background on Prior-data Fitted Networks and TabPFN
Prior-data Fitted Networks (PFNs) are a class of models that learn to approximate Bayesian inference for a given prior (Müller et al., 2022). Instead of fitting a new model from scratch, a PFN is an individual, large pre-trained Transformer to perform classification or regression in a single forward pass. This process, known as in-context learning (ICL), allows the model to make predictions using sequences of labeled examples provided directly in the input, without requiring any gradient updates (Dong et al., 2024). The entire prediction algorithm is contained in the weights of the network, which is trained once on millions of synthetically generated datasets sampled from the prior. At inference time, the trained PFN takes a real-world dataset, composed of training and test samples, as a set-valued input and returns a distribution over the output space. This output space is categorical for classification tasks and the real line for regression tasks.
Posterior predictive modeling and synthetic prior fitting.
PFNs are rooted in Bayesian supervised learning, where the primary objective is to model the posterior predictive distribution (PPD) (MacKay, 1992; Seeger, 2004; Blei et al., 2017). Since computing the PPD is often intractable (MacKay, 1992), PFNs instead learn to approximate the PPD offline through a process called synthetic prior fitting Müller et al. (2022). This is achieved using a prior specified by a sampling scheme that first samples a data-generating mechanism, $\phi\sim p(\phi)$, and then samples a synthetic dataset, $D\sim p(D|\phi)$. This process is repeated to generate millions of diverse datasets for training. The network’s parameters, $\theta$, are then optimized to predict held-out test samples ($D_{\text{test}}\subset D$) conditioned on the rest of the dataset ($D_{\text{train}}=D\setminus D_{\text{test}}$). The training objective is to minimize the negative log likelihood (NLL) loss on these held-out examples, averaged over the prior distribution of datasets $p(D)$:
$\mathcal{L}_{\text{NLL}}(\theta)=\mathbb{E}_{((x_{\text{test}},y_{\text{test}})\cup D_{\text{train}})\sim p(D)}[-\log q_{\theta}(y_{\text{test}}|x_{\text{test}},D_{\text{train}})].$ (2)
This training process makes it explicit that PFNs are trained to emulate Bayesian inference by averaging over tasks drawn from a prior distribution over data-generating functions. Minimizing this loss ensures that the trained neural network, $q_{\theta}$, learns to approximate the true Bayesian PPD for the specified prior.
TabPFN.
TabPFN is a PFN built specifically for tabular supervised learning (Hollmann et al., 2023). The model employs a novel two-way attention mechanism specifically designed for tabular data. Unlike standard transformers (Vaswani et al., 2023) that treat tabular data as sequential tokens, TabPFN assigns separate representations to each cell in the table. The architecture uses alternating attention patterns: each cell first attends to other features within its row (inter-feature attention), then attends to the same feature across all rows (inter-sample attention). This design ensures permutation invariance for both samples and features while enabling efficient scaling to larger tables than those seen during training. TabPFN v2 (Hollmann et al., 2025) retains the core training paradigm of the original TabPFN while introducing several key enhancements that improve accuracy, runtime, and applicability. Going forward, when referring to TabPFN, we mean TabPFN v2 since it is the most up-to-date version.
3 Training TabImpute on Synthetic Data
We develop a featurization for tabular missing data imputation that enables better utilization of TabPFN’s architecture, create a synthetic data generation pipeline across diverse missingness patterns, and employ an adaptive training algorithm to balance performance across all patterns. Training used $8$ H200 GPUs and an Intel Xeon Platinum 8592+ CPU over approximately one week, processing $25$ million synthetic tables. Our model matches TabPFN’s size and runs on CPU-only systems. Evaluation used $1$ H200 GPU.
3.1 Entry-wise Featurization and Architecture
We recast missing data imputation as supervised learning to leverage TabPFN’s architecture and enable parallel GPU computation of missing entries. For each dataset (training point), let $X^{*}$ be the complete matrix with $m$ rows and $n$ columns, $\Omega$ be the set of missing entry indices, and $X$ be the matrix with induced missingness:
$X_{ij}=\begin{cases}X^{*}_{ij}&\text{for}\quad(i,j)\in[m]\times[n]\setminus\Omega\\ \star&\text{otherwise.}\end{cases}$ (3)
where $\star$ denotes a missing entry. Let $\Omega_{\text{obs}}=[m]\times[n]\setminus\Omega$. Our feature matrix construction adds $(i\oplus j\oplus X_{i,:}\oplus X_{:,j})$ for each entry $i,j\in[m]\times[n]$, where $X_{i,:}$ denotes the $i$-th row, $X_{:,j}$ the $j$-th column, and $\oplus$ concatenation. Each row’s target is $y_{ij}=X^{*}_{ij}$. During pre-training, we train the model to predict target values for all $(i,j)\in\Omega$. This procedure creates a feature matrix of size $nm\times(n+m)$. This featurization captures all necessary information for each cell through its row and column context while enabling parallel computation of missing entries on GPUs. Although the input matrix size increases, parallelization gains outweigh this cost.
Architecture.
We use TabPFN’s base architecture with one modification: removing the attention mask to allow training points to attend to test points. Since our train/test set sizes vary randomly with missingness patterns (unlike TabPFN’s controlled synthetic generation), we remove the mask to enable parallel batch training. TabPFN’s mask prevents train points from seeing test feature distributions, which is important in general supervised learning. However, our test set is created using data already available to the observed points, thus alleviating any data-leakage concerns.
3.2 Synthetic Training Data Generation
We generate tens of millions of datasets with missing values to train our model through a two-step process: first, generating underlying data, then introducing missingness patterns on top.
3.2.1 Data Generation with Linear Factor Models
We generate data using a simple linear factor model (LFMs) (Bai & Ng, 2002). LFMs are commonly used in matrix completion literature to prove error bounds for matrix completion algorithms (Koren et al., 2009; Candes & Recht, 2012). This family of models generates a data matrix $Y\in\mathbb{R}^{m\times n}$ by assuming the data lies on or near a low-dimensional subspace. The simplest case generates the data matrix $Y$ as the inner product of two lower-rank latent factor matrices, $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$, where $k\ll n,m$ is the rank:
$Y=UV^{T}.$ (4)
To generate diverse datasets, the latent vectors (rows of $U$ and $V$) are sampled from a variety of distributions, including Gaussian, Laplace, Student’s t, spike-and-slab (a mixture of a Dirac delta at zero and a Gaussian), and Dirichlet.
When training, we experimented with several classes of data-generating processes, including matrices from nonlinear factor models and structural causal models (SCM) similar to the ones used in TabPFN. However, we found that the model trained on linear factor models performed the best. See Tab. 5 in Sec. A.3 for results demonstrating this.
3.2.2 Missingness Patterns for Training and Evaluation
After generating a complete data matrix $X^{*}\in\mathbb{R}^{m\times n}$, we introduce missingness by applying a masking matrix $M\in\{0,1\}^{m\times n}$, where the entry value $M_{ij}=1$ if and only if $(i,j)\in\Omega_{\mathrm{obs}}$. To ensure TabImpute is robust and generalizable to the variety of ways data can be missing in real-world scenarios, we pre-train on a comprehensive stock of synthetic datasets with several missingness patterns. For convenience, we define $p_{ij}=\mathbb{P}(M_{ij}=1)$, the propensity of each entry in $X^{*}$.
We include $13$ different missingness patterns: $1$ MCAR, $1$ MAR pattern, and $11$ MNAR patterns. For examples of these, see Fig. 2. MNAR patterns often stump standard imputation methods and yet are extremely common in the real world.
MCAR:
MCAR missingness means the probability of an entry being missing, defined as its propensity, is constant across entries and independent from any other randomness. The missingness indicators $M_{ij}$ are drawn i.i.d. from a Bernoulli distribution $M_{ij}\sim\mathrm{Bern}(p)$ across all $(i,j)\in[m]\times[n]$ for some constant $p\in(0,1)$. This is the simplest form of missingness, but is unrealistic (Van Buuren, 2012).
MAR:
For MAR missingness, the probability of an entry being missing depends only on the observed values $X$. In other words, the randomness in MAR can be explained by conditioning on observed factors. Additionally, every entry has a positive probability of being observed (i.e., $p_{ij}>0$). We simulate MAR through column-wise MAR, denoted Col-MAR: we choose several columns as predictor columns and use those values to mask entries in other columns. This is similar to the MAR approach taken in Jarrett et al. (2022).
MNAR Patterns:
For the most complex missingness class, MNAR, the probability of an entry being missing can depend on unobserved factors. Note that MNAR patterns are significantly more difficult to handle systematically, often requiring specialized methods for a specific kind of MNAR pattern (Van Buuren, 2012). Due to the flexibility of our entry-wise featurization, TabImpute can produce imputations for these highly complex scenarios, even when columns are completely missing, such as in panel-data missingness patterns. HyperImpute was tested on two MNAR patterns in the Appendix of Jarrett et al. (2022), one where values are further masked after an MAR pattern and another where values outside a certain range are masked. We build on this work by testing on $11$ MNAR patterns (see Sec. A.2 and Tab. 6 for details). We implement a range of MNAR patterns to simulate plausible real-world scenarios. For example, we utilize the expressiveness of neural networks to create random propensity functions MNAR patterns, use bandit algorithms to induce column-adaptive missing patterns, simulate panel data missingness where some features are removed later, censoring where sensor readings fall outside a detectable range, and survey data artifacts like respondent polarization and skip-logic.
3.3 Training on Multiple Missingness Patterns
We train our model to predict unobserved values under several missingness patterns simultaneously. For the final TabImpute model, we trained only on MCAR, NN-MNAR, and Self-Masking-MNAR because we found our model generalized to the other methods well without including them explicitly in training. We use the prior-data fitted negative log likelihood (NLL) loss proposed in Müller et al. (2022). Like other PFNs with continuous numerical output, we use the Riemann distribution output also proposed in Müller et al. (2022). Since we can generate an unlimited amount of synthetic data, we do not reuse any synthetic data and only do one gradient pass per batch of datasets. This allows our model to learn the underlying data-generating process and missingness mechanisms without risk of memorization. We use a learning rate of $0.0001$, a batch size of $64$, and train on around $25$ million synthetic datasets.
We use an adaptive algorithm to determine what proportion of missingness patterns we include in each batch as we train: Every $s$ gradient steps, we recalculate the proportion of each missingness pattern by: (i) create new batches of each missingness pattern, (ii) calculate the model’s loss value for each missingness pattern’s batch, and (iii) recalculate proportions by running softmax (Heaton, 2018, Pgs. 180-184) over the loss values. This procedure adaptively down-weights any missingness type our model performs well on and up-weights any missingness type our model performs poorly on. For our pre-trained TabImpute model, we set $s=50$ steps.
Remark: We had initially attempted to train the model sequentially on one missingness pattern at a time, but found that the network always experienced catastrophic forgetting (McCloskey & Cohen, 1989) irrespective of learning rate (i.e., it forgot how to handle the previous missingness patterns).
3.4 Adaptive Ensembling
Our final method, TabImpute+, includes two layers of ensembling: (i) permutation-based pre-processing techniques and (ii) adaptive weighted average of TabImpute and EWF-TabPFN. Note that TabPFN’s regression class includes several ensembling techniques as well. The first level of ensembling performs random permutations of the rows and columns before running these permuted matrices through TabImpute, reversing the permutations, and then averaging the predictions across the runs. We used $4$ independent permutations for our final method.
[…]
4 Results on OpenML Datasets: MissBench
To evaluate TabImpute against other methods, we introduce MissBench: a missing-data imputation benchmark using $42$ OpenML (Vanschoren et al., 2013) tabular datasets with $13$ synthetic missingness patterns. For every dataset and missingness pattern, we test each method’s ability to impute masked values. The $42$ OpenML datasets span domains such as medicine, engineering, and education. The missingness patterns include $1$ MCAR pattern, $1$ MAR pattern, and $11$ MNAR patterns. We provide details for the MNAR patterns in Sec. A.2.
We test a suite of imputation methods on MissBench: column-mean imputation (Hawthorne & Elliott, 2005), SoftImpute (Hastie et al., 2015), MissForest (Stekhoven & Bühlmann, 2011), iterative chained estimators (ICE/MICE) (van Buuren & Groothuis-Oudshoorn, 2011; Royston & White, 2011), GAIN (Yoon et al., 2018), MIWAE (Mattei & Frellsen, 2019), an optimal transport-based method (Muzellec et al., 2020), and TabPFN’s imputation method (see Sec. 1).
Imputation Accuracy.
To ensure a fair comparison across datasets with different scales and inherent difficulties, we report a normalized accuracy score. In particular, for each dataset and missingness pattern, we first calculate the standard Root Mean Squared Error (RMSE) for every imputation method as $\big(\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}\big(X^{\text{true}}_{ij}-X^{\text{imputed}}_{ij}\big)^{2}\big)^{1/2}$}, where $\Omega$ denotes the set of missing entries. We then perform a min-max normalization on these RMSE scores across all methods for that specific task:
$\text{Normalized RMSE}=\frac{\text{RMSE}_{\text{method}}-\min(\text{RMSE}_{\text{all methods}})}{\text{max}(\text{RMSE}_{\text{all methods}})-\min(\text{RMSE}_{\text{all methods}})}$ (5)
This normalization maps the best-performing method to 0 and the worst to 1. Finally, we define our Imputation Accuracy as 1Normalized RMSE1-$\text${Normalized RMSE}, where higher values indicate better performance.
The final scores we report represent imputation accuracy averaged across all $42$ datasets and $13$ missingness patterns. Dataset sizes range from $50\times 5$ to $170\times 55$. We evaluate only on datasets with numerical values without pre-existing missingness before applying synthetic patterns. Specific datasets are listed in Tab. 7 (Sec. A.3).
Tab. 1 presents results for each missingness pattern as well as overall performance. TabImpute+ achieves the best overall performance and for nearly all individual patterns. For completeness, we list the performance of methods not shown in the table in Sec. A.3 with non-normalized RMSE examples in Tab. 8. TabImpute+ performs best under high missingness conditions (Fig. 4, Sec. A.1), which is expected since it leverages generative pre-training rather than relying solely on available dataset information like discriminative methods.
Table 1: Imputation Accuracy ± Standard Deviation by Missingness Pattern.
We train on the patterns above the dashed line.
Pattern TabImpute+ HyperImpute OT MissForest
MCAR 0.821 ± 0.157 0.804 ± 0.205 0.856 ± 0.114 0.867 ± 0.148
NN-MNAR 0.880 ± 0.126 0.762 ± 0.240 0.815 ± 0.139 0.819 ± 0.188
Self-Masking-MNAR 0.707 ± 0.279 0.717 ± 0.257 0.593 ± 0.246 0.666 ± 0.272
[2pt/2pt] Col-MAR 0.860 ± 0.178 0.819 ± 0.263 0.756 ± 0.262 0.776 ± 0.251
Block-MNAR 0.908 ± 0.168 0.873 ± 0.178 0.857 ± 0.152 0.860 ± 0.153
Seq-MNAR 0.905 ± 0.094 0.862 ± 0.193 0.892 ± 0.094 0.829 ± 0.203
Panel-MNAR 0.791 ± 0.329 0.515 ± 0.368 0.540 ± 0.374 0.526 ± 0.371
Polarization-MNAR 0.885 ± 0.137 0.620 ± 0.263 0.804 ± 0.125 0.559 ± 0.231
Soft-Polarization-MNAR 0.864 ± 0.176 0.705 ± 0.223 0.755 ± 0.195 0.670 ± 0.269
Latent-Factor-MNAR 0.887 ± 0.121 0.778 ± 0.231 0.844 ± 0.167 0.836 ± 0.152
Cluster-MNAR 0.873 ± 0.135 0.841 ± 0.183 0.828 ± 0.137 0.833 ± 0.141
Two-Phase-MNAR 0.855 ± 0.186 0.871 ± 0.200 0.809 ± 0.183 0.878 ± 0.123
Censoring-MNAR 0.594 ± 0.302 0.797 ± 0.262 0.599 ± 0.239 0.682 ± 0.249
Overall 0.833 ± 0.213 0.766 ± 0.259 0.765 ± 0.227 0.754 ± 0.248
Building on TabPFN.
The TabPFN team’s imputation extension predicts missing entries column-by-column using their pre-trained model. However, as shown in Fig. 1, this method has poor accuracy and slow runtime even on H200 GPU’s. This motivated us to introduce a new entry-wise featurization that achieves improved performance using TabPFN’s model, which we denote as EWF-TabPFN. We initially attempted fine-tuning TabPFN with our synthetic data but consistently encountered catastrophic forgetting, with performance degrading to match TabImpute, the model trained from scratch. Without access to TabPFN’s original training data, proper data augmentation proved immensely challenging. Instead, we elected to ensemble TabImpute and EWF-TabPFN as TabImpute+, leveraging their complementary performance across benchmark patterns (Tab. 2). Crucially, the performance scores in this table are re-normalized using only the results of these three methods. This means the minimum and maximum RMSE values used for normalization are determined from this smaller set, leading to different numerical scores than in Tab. 1, which was normalized across all 12 competing methods. This approach offers a clearer view of the relative performance and complementarity within our proposed family of models, highlighting which components excel at different missingness patterns. (See Sec. A.3 for other test results.)
Table 2: Imputation Accuracy ± Standard Deviation for Zero-Shot Methods.
Pattern TabImpute+ EWF-TabPFN TabImpute
MCAR 0.349 ± 0.468 0.324 ± 0.465 0.667 ± 0.477
NN-MNAR 0.551 ± 0.480 0.458 ± 0.488 0.487 ± 0.488
Self-Masking-MNAR 0.945 ± 0.214 0.952 ± 0.211 0.048 ± 0.216
[2pt/2pt] Col-MAR 0.799 ± 0.391 0.773 ± 0.414 0.220 ± 0.419
Block-MNAR 0.498 ± 0.474 0.437 ± 0.494 0.559 ± 0.501
Seq-MNAR 0.394 ± 0.482 0.363 ± 0.447 0.619 ± 0.492
Panel-MNAR 0.565 ± 0.476 0.568 ± 0.476 0.415 ± 0.499
Polarization-MNAR 0.775 ± 0.407 0.788 ± 0.411 0.214 ± 0.415
Soft-Polarization-MNAR 0.698 ± 0.422 0.791 ± 0.383 0.258 ± 0.439
Latent-Factor-MNAR 0.505 ± 0.490 0.467 ± 0.480 0.524 ± 0.505
Cluster-MNAR 0.428 ± 0.474 0.420 ± 0.487 0.571 ± 0.501
Two-Phase-MNAR 0.591 ± 0.475 0.560 ± 0.494 0.429 ± 0.501
Censoring-MNAR 0.891 ± 0.304 0.902 ± 0.282 0.098 ± 0.300
Overall 0.614 ± 0.468 0.600 ± 0.476 0.393 ± 0.487
5 Conclusion & Future Work
In this paper, we introduce a comprehensive benchmark for tabular missing data, MissBench, and a pre-trained transformer for the tabular missing data problem, TabImpute. We build on recent work in tabular representation learning by adapting TabPFN’s architecture and training pipeline for the missing data setting. While we train purely on synthetic data, we are able to impute entries accurately on real-world OpenML data for a comprehensive set of missingness patterns, showcasing our model’s ability to generalize to unseen domains. We open-source our model architecture, weights, and our training and evaluation code (available at https://github.com/jacobf18/tabular). We hope this will facilitate others to utilize, validate, and build on our work.
Since we use the same architecture as TabPFN, we also suffer from the quadratic time complexity of attention along both rows and columns. Due to entry-wise featurization, this complexity is squared again along the row axis. While TabImpute proved to be fast on relatively small tables in MissBench, we expect scalability to be a concern on larger table sizes. Additionally, TabImpute is very fast when using a GPU, but is slower on a CPU similar to TabPFN. This can be partly alleviated by imputing tables in chunks, but we leave this for future work.
Also note that since TabImpute uses the same architecture as TabPFN, any improvements to TabPFN’s architecture can be immediately ported to TabImpute. For example, Zeng et al. (2025) and Qu et al. (2025) propose different attention mechanisms to speed up TabPFN-like architectures for tabular data. Finally, since we use a PFN architecture, we output a distribution for each missing entry and can sample from this distribution for multiple imputation (Rubin, 2018). In future work, we plan on (i) exploring further training on more complex missingness patterns and data-generating processes, (ii) enhancing our method to support categorical data, (iii) extending our evaluation set to causal inference settings, which can be modeled as missing-data problems (Agarwal et al., 2023), (iv) improving the architecture to scale to larger datasets, and (v) utilizing our method for multiple imputation.
Appendix A Appendix
Here we present the rest of the missingness patterns we tested on, tables with further results, the methods we tested against, and the OpenML datasets we evaluated on.
A.1 Additional tests
Next, we discuss when TabImpute+ does well and when HyperImpute and other methods do well. We found that an important factor in determining performance was the level of missingness. The probability of missingness can only be controlled precisely for MCAR. Thus, we show in Fig. 4 the performance of the top methods as we increase the missingness level.
Figure 4: Imputation accuracy versus fraction of missingness for MCAR. TabImpute+ performs the best when missingness is higher because it is a generative model that fits to the data in context.
A.2 Details for MNAR Missingness Patterns
This section provides the mathematical and implementation details for each of the simulated MNAR missingness mechanisms that we implement and test. For each pattern, we define the mechanism by which the missingness mask $M$ is generated, where $M_{ij}=1$ if the value $X_{ij}$ is observed and $M_{ij}=0$ otherwise.
A.2.1 Details for NN-MNAR
Description
The pattern simulates a scenario where the propensity $p_{ij}$ depends on the underlying matrix values $X^{*}$ in an arbitrary manner. We achieve a comprehensive coverage of MNAR patterns by leveraging the expressiveness of neural networks.
Methodology
One general form of MNAR can be described as follows: for all $i$ and $j$, there exists some function $f_{ij}$ on the true (hence unobserved) matrix $X^{*}$ such that the propensity depends on $X^{*}$ as follows: $p_{ij}(X^{*})=\mathbb{P}(M_{ij}=1|X^{*})=f_{ij}(X^{*})$. By leveraging the expressiveness of neural networks, we propose a neural-net-based MAR pattern generator (NN-MNAR) that is designed to approximate arbitrary propensities characterized by functions $f_{ij}$.
Implementation Details
For fixed indices $i$ and $j$, NN-MNARconstructs the propensity $p_{ij}$ in a two-step procedure. First, we randomly collect a subset of values from the matrix $X^{*}$ and flatten them as a vector, say $X^{*}(i,j)$; we do this by first randomly generating a neighborhood $\mathbf{N}_{ij}\subset[m]\times[n]$, then the entries in the neighborhood $X^{*}_{st},(s,t)\in\mathbf{N}_{ij}$ constitute the entries of the vector $X^{*}(i,j)$. Second, a neural-net function $g_{ij}:\mathbb{R}^{|\mathbf{N}_{ij}|}\to[0,1]$ is constructed by randomly initializing the number of layers, depth, weight, and bias. At each training step, the random neighborhood $\mathbf{N}_{ij}$ and the random neural-net $g_{ij}$ collectively defines the propensity $p_{ij}=g_{ij}(X^{*}(i,j))$ from which MNAR missingness patterns are generated $M_{ij}\sim\mathrm{Bern}(p_{ij})$.
A.2.2 Details on Seq-MNAR
Description
This pattern simulates a scenario where masking matrix values $M_{ij}\in\{0,1\}$ for each column $j$ are adaptively chosen depending on the information up to column j1j-1 (i.e., regard columns as time). Specifically, we employ variants of bandit algorithms Lattimore & Szepesvári (2020) while regarding the binary masking matrix values as the two arms. Such patterns commonly arise in sequential experiments Ghosh et al. (2024).
Methodology
The true matrix $X^{*}$ is transformed to constitute the reward. For each designated column $j$, one of the following bandit algorithm utilizes the all the information of $X^{*}$ and $M$ up to column j1j-1 and chooses one of the two arms $\{0,1\}$ via one of the following algorithms: $\varepsilon$-greedy, upper-confidence bound (UCB), Thompson-sampling (Thompson, 1933) or gradient bandit.
Implementation Details
We generate exogenous Gaussian noise and add it to the true matrix $X^{*}$ and regard $X^{*}$ as the reward for arm 0 and its noisy version as the reward for arm $1$. Then, starting from the first column with multiple rows as multiple agents, we randomly initiate (with random configurations) one of the four algorithms Lattimore & Szepesvári (2020): $\varepsilon$-greedy, Upper Confidence Bound (UCB) (Auer et al., 2002), Thompson sampling with random configurations (Thompson, 1933). Further, we have the option to randomly mix pooling techniques Ghosh et al. (2024) on top of any of the four algorithms.
A.2.3 Self-Masking-MNAR
Description:
This pattern simulates a scenario where the probability of a value being missing is a direct function of the value itself. This pattern can be regarded as a special case of NN-MNAR. A common real-world example includes individuals with very high incomes being less likely to report their salary on a survey.
Methodology:
For each designated target column $j$, the probability of an entry $(i,j)$ being missing is determined by a logistic function of its value. The relationship is defined as $\mathbb{P}(M_{ij}=0|X^{*}_{ij})=\sigma(\alpha\cdot X^{*}_{ij}+\beta_{0})$ where $\sigma(z)=(1+e^{-z})^{-1}$ is the sigmoid function.
Implementation Details:
A random coefficient $\alpha$ is chosen from the set $\{-2,-1,1,2\}$ to introduce variability in the direction and magnitude of the value’s effect on its missingness probability. The bias term $\beta_{0}$ is calibrated to achieve a target missingness proportion, $p$.
A.2.4 Censoring-MNAR
Description:
This pattern models missingness that arises from the practical/physical limits of measurement equipment, where values below a lower detection limit (left-censoring) or above an upper detection limit (right-censoring) are not recorded. We would expect to see such missingness in sensor or biological assay datasets where test equipment can only detect biomarker levels above/below certain thresholds.
Methodology:
For each column $j$, a censoring direction (left or right) is chosen with equal probability. A cutoff value is determined based on a specified quantile, $q_{\text{censor}}$, of the set of currently observed (non-missing) values in that column. Let $X^{*}_{:,j}$ denote the set of observed values in column $j$, i.e., $X^{*}_{:,j}=\{X_{ij}\mid M_{ij}=1\}$.
  • Left-Censoring: All values in column $j$ that are less than the $q_{\text{censor}}$-th quantile of the observed values in that same column are set to missing. The threshold is a single scalar value calculated from the column’s observed data.
    $M_{ij}=0\quad\text{if}\quad X_{ij}<\text{quantile}(X^{*}_{:,j},q_{\text{censor}})$
  • Right-Censoring: All values in column $j$ that are greater than the $(1-q_{\text{censor}})$-th quantile of the observed values in that column are set to missing. The threshold is a single scalar value.
    $M_{ij}=0\quad\text{if}\quad X_{ij}>\text{quantile}(X^{*}_{:,j},1-q_{\text{censor}})$
Implementation Details:
The choice between left- and right-censoring is made randomly for each column with a probability of 0.5 for each. We introduce a hyperparameter $q_{\text{censor}}$ for the censoring quantile that controls the fraction of data to be censored from either tail of the distribution. For our evaluation, we use $q_{\text{censor}}=0.25$.
A.2.5 Panel-MNAR
Description:
This pattern simulates participant dropout in longitudinal or panel data studies, where once a subject (row) drops out at a specific time point, all their subsequent data is missing. This is a common occurrence in clinical trials or long-term studies/surveys.
Methodology:
The columns of the data matrix $X^{*}$ are assumed to represent ordered time points t=0,1,,T1t=0,1,...,T-1. For each subject (row) $i$, a random dropout time $t_{0,i}$ is sampled. All observations for that subject from time $t_{0,i}$ onwards are masked as missing.
$M_{ij}=0\quad\forall j\geq t_{0,i}$
Implementation Details:
For each row $i$, the dropout time $t_{0,i}$ is sampled uniformly from the range of possible time steps, i.e., $t_{0,i}\sim\mathrm{Unif}\{1,...,T-1\}$.
A.2.6 Polarization-MNAR
Description:
Values falling in the middle of a feature’s distribution are preferentially removed, simulating survey non-response from individuals with moderate opinions. This is implemented by setting values between the $q$-th and (1q)(1-q)-th quantiles to missing. A ”soft” version makes the observation probability proportional to the value’s distance from the median. Such patterns are most commonly seen in political polls, where moderate voters are less likely to respond while those with extreme views are more likely to respond, or in product reviews, where only very satisfied or very dissatisfied customers leave ratings.
Hard Polarization Methodology:
For each column $j$, values falling between two quantiles are deterministically masked. The quantiles are calculated using only the set of currently observed (non-’NaN’) values in that column. Let $X^{*}_{:,j}$ denote this set of observed values, i.e., $X^{*}_{:,j}=\{X_{ij}\mid M_{ij}=1\}$. The lower and upper thresholds, $L_{j}$ and $H_{j}$, are defined as:
$L_{j}=\text{quantile}(X^{*}_{:,j},q_{\text{thresh}})$
$H_{j}=\text{quantile}(X^{*}_{:,j},1-q_{\text{thresh}})$
An entry $X_{ij}$ is then masked if its value falls between these two scalar thresholds:
$M_{ij}=0\quad\text{if}\quad L_{j}<X_{ij}<H_{j}$
Soft Polarization Methodology:
The probability of a value being observed is made proportional to its normalized absolute distance from the column’s median, $\mu_{j}$. This creates a softer, probabilistic version of the polarization effect. The missing probability is given by:
$\mathbb{P}(M_{ij}=0)=\epsilon+(1-2\epsilon)\frac{|X^{*}_{ij}-\mu_{j}|^{\alpha}}{\max_{k}(|X^{*}_{kj}-\mu_{j}|^{\alpha})}$
Implementation Details:
For the hard polarization pattern, we introduce a hyperparameter $\bm{q_{\text{thresh}}}$ for the threshold quantile that defines the central portion of the distribution to be masked. For the soft polarization pattern, we have an exponent parameter $\alpha$ that controls the intensity of the polarization effect. Higher values of $\alpha$ make the observation probability more sensitive to deviations from the median. In the soft version, we also have a baseline probability $\epsilon$ that ensures even values at the median have a non-zero chance of being observed.
A.2.7 Latent-Factor-MNAR
Description:
This pattern generates a complex missingness structure where the probability of an entry being missing depends on unobserved (latent) characteristics of both its row and its column. This is common in recommender systems, where a user’s decision to rate an item depends on latent user preferences and item attributes. Other real-world examples include job application screenings where certain combinations of demographic factors influence whether candidates complete applications, independent of their qualifications, or online dating profiles where cultural background affects disclosure of personal information.
Methodology:
The probability of an entry $(i,j)$ being observed is modeled using a low-rank bilinear model. The observation probability is given by the sigmoid of a dot product of latent factors plus bias terms:
$\mathbb{P}(M_{ij}=1)=\sigma(u_{i}^{T}v_{j}+b_{i}+c_{j})$
where $u_{i}\in\mathbb{R}^{k}$ and $v_{j}\in\mathbb{R}^{k}$ are $k$-dimensional latent vectors for row $i$ and column $j$, and $b_{i}$ and $c_{j}$ are scalar biases for the row and column, respectively.
Implementation Details:
We specify the rank $k$ that defines the dimensionality of the latent space, sampled as an integer. The elements of the latent factor matrices $U\in\mathbb{R}^{N\times k}$ and $V\in\mathbb{R}^{D\times k}$, and the bias vectors $b\in\mathbb{R}^{N}$ and $c\in\mathbb{R}^{D}$, are sampled independently from a standard normal distributions.
A.2.8 Cluster-MNAR
Description:
This pattern induces missingness based on latent group-level characteristics. Rows and columns are first assigned to discrete clusters, and each cluster has a random effect that uniformly influences the observation probability of all its members. This is useful for modeling data from grouped experiments, such as a clinical trial where patients (rows) from a specific hospital (a given row cluster) and certain lab tests (a column cluster) might have systematically different rates of missingness due to local care protocols or equipment availability.
Methodology:
The probability of an entry $(i,j)$ being observed is determined by an additive model of random effects corresponding to the cluster assignments of its row $i$ and column $j$. Denoting the row assignments by $C_{R}(i)$ and column assignments by $C_{C}(j)$, the observation probability is modeled as:
$\mathbb{P}(M_{ij}=1)=\sigma(g_{C_{R}(i)}+h_{C_{C}(j)}+\epsilon_{ij})$
where:
  • $\sigma(z)=(1+e^{-z})^{-1}$ is the sigmoid function.
  • $g_{k}\sim\mathcal{N}(0,\tau_{r}^{2})$ is the random effect for row cluster $k$.
  • $h_{l}\sim\mathcal{N}(0,\tau_{c}^{2})$ is the random effect for column cluster $l$.
  • $\epsilon_{ij}\sim\mathcal{N}(0,\epsilon_{\text{std}}^{2})$ is an entry-specific noise term.
Implementation Details:
For a matrix with $N$ and $D$ columns, row assignments $C_{R}(i)$ are draw uniformly from $\{0,...,K_{R}-1\}$ and column assignments $C_{C}(j)$ are drawn uniformly from $\{0,...,K_{C}-1\}$, where $K_{R}$ and $K_{C}$ are the total number of row and column clusters, respectively. The number of row clusters $K_{R}$, the number of column clusters $K_{C}$, and the standard deviation of the random effects ($\tau_{r}$, $\tau_{c}$, $\epsilon_{\text{std}})$ are hyperparameters specified for the data generation process.
A.2.9 Two-Phase-MNAR
Description:
This mechanism mimics multi-stage data collection where a subset of participants from an initial survey (with “cheap” features) are selected for a more detailed follow-up survey (with “expensive” features). The selection for the second phase is dependent on the data collected in the first, making this an MNAR pattern. A specific example includes market research where basic demographics are collected from all participants, but detailed purchasing behavior is only gathered from a subset, with missingness related to income level.
Methodology:
Let $\mathcal{F}=\{0,1,...,D-1\}$ be the set of all column indices in the data matrix $X$. This set is randomly partitioned into a ”cheap” subset $\mathcal{C}\subset\mathcal{F}$ and an ”expensive” subset $\mathcal{E}\subset\mathcal{F}$, such that $\mathcal{C}\cup\mathcal{E}=\mathcal{F}$ and $\mathcal{C}\cap\mathcal{E}=\emptyset$. By design, features in the cheap set $\mathcal{C}$ are always observed.
The decision to collect the expensive features for a given row $i$ is based on a logistic model applied to its cheap features. Let $X_{i,\mathcal{C}}$ denote the vector of values $\{X_{ij}\mid j\in\mathcal{C}\}$ for row $i$. A score is calculated for each row:
$s_{i}=\text{normalize}(X_{i,\mathcal{C}}^{T}w)$
where $w$ is a vector of random weights and the ‘normalize‘ function applies z-score normalization to the resulting scores across all rows.
The probability that all expensive features are observed for row $i$ is then given by:
$\mathbb{P}(M_{ij}=1\text{ for all }j\in\mathcal{E})=\sigma(\alpha+\beta\cdot s_{i})$
If the expensive features for row $i$ are not observed (based on the probability above), then all of its values in the expensive columns are masked as missing, i.e., $M_{ij}=1$ for all $j\in\mathcal{E}$.
Implementation Details:
A fraction of columns, e.g., 50%, are randomly assigned to be ”cheap”. The weight vector for the scoring model is sampled from a standard normal distribution, $w\sim\mathcal{N}(0,1)$. Parameters $\alpha,\beta$ control the base rate and score-dependency of the observation probability. In our implementation, they are set to default values of $\alpha=0$ and $\beta=2.0$.
A.3 Additional tables
Table 3: Other imputation methods
Name Description
Column-wise mean (Hawthorne & Elliott, 2005) Mean of columns
SoftImpute (Hastie et al., 2015) Iterative soft thresholding singular value decomposition based on a low-rank assumption on the data
$k$-Nearest Neighbors (Fix & Hodges, 1989) Row-wise nearest neighbors mean
HyperImpute (Jarrett et al., 2022) Iterative imputation method optimizing over a suite of imputation methods
Optimal transport method (Muzellec et al., 2020) Uses optimal transport distances as a loss to impute missing values based on the principle that two randomly drawn batches from the same dataset should share similar data distributions
MissForest (Stekhoven & Bühlmann, 2011) Repeatedly trains a random forest model for each variable on the observed values to predict and fill in missing entries until convergence
ICE (van Buuren & Groothuis-Oudshoorn, 2011) Imputation with iterative and chained equations of linear/logistic models for conditional expectations
MICE (Royston & White, 2011) Handles missing data by iteratively imputing each incomplete variable using regression models that condition on all other variables
GAIN (Yoon et al., 2018) Adapts generative adversarial networks (Goodfellow et al., 2020) where the generator imputes missing values and the discriminator identifies which components are observed versus imputed
MIWAE (Mattei & Frellsen, 2019) Learns a deep latent variable model and then performs importance sampling for imputation
ForestDiffusion (Jolicoeur-Martineau et al., 2024) Trains a diffusion model using XGBoost directly on incomplete tabular data and then fills in missing values with an adapted inpainting algorithm
Table 4: Imputation Accuracy ± Standard Deviation by Missingness Pattern (Remaining Methods)
Pattern K-Nearest Neighbors ICE Col Mean TabPFN
MCAR 0.762 ± 0.158 0.658 ± 0.273 0.473 ± 0.263 0.410 ± 0.257
Col-MAR 0.846 ± 0.183 0.832 ± 0.248 0.627 ± 0.299 0.553 ± 0.330
NN-MNAR 0.757 ± 0.168 0.643 ± 0.254 0.507 ± 0.266 0.437 ± 0.266
Block-MNAR 0.857 ± 0.184 0.813 ± 0.239 0.774 ± 0.241 0.572 ± 0.335
Seq-MNAR 0.889 ± 0.114 0.819 ± 0.240 0.778 ± 0.214 0.641 ± 0.290
Self-Masking-MNAR 0.623 ± 0.268 0.681 ± 0.328 0.292 ± 0.258 0.268 ± 0.248
Panel-MNAR 0.614 ± 0.400 0.584 ± 0.407 0.502 ± 0.389 0.442 ± 0.384
Polarization-MNAR 0.631 ± 0.199 0.547 ± 0.279 0.972 ± 0.059 0.914 ± 0.187
Soft-Polarization-MNAR 0.575 ± 0.266 0.723 ± 0.217 0.754 ± 0.207 0.598 ± 0.327
Latent-Factor-MNAR 0.776 ± 0.186 0.716 ± 0.268 0.601 ± 0.297 0.508 ± 0.316
Cluster-MNAR 0.784 ± 0.150 0.758 ± 0.279 0.559 ± 0.278 0.500 ± 0.297
Two-Phase-MNAR 0.904 ± 0.124 0.881 ± 0.208 0.666 ± 0.273 0.466 ± 0.329
Overall 0.751 ± 0.238 0.721 ± 0.289 0.625 ± 0.313 0.526 ± 0.334
Pattern MIWAE MICE GAIN
MCAR 0.236 ± 0.274 0.316 ± 0.283 0.543 ± 0.256
Col-MAR 0.491 ± 0.315 0.521 ± 0.354 0.276 ± 0.354
NN-MNAR 0.234 ± 0.280 0.310 ± 0.325 0.400 ± 0.324
Block-MNAR 0.594 ± 0.267 0.491 ± 0.311 0.184 ± 0.317
Seq-MNAR 0.586 ± 0.289 0.478 ± 0.329 0.204 ± 0.311
Self-Masking-MNAR 0.209 ± 0.269 0.631 ± 0.348 0.456 ± 0.380
Panel-MNAR 0.374 ± 0.363 0.251 ± 0.325 0.279 ± 0.360
Polarization-MNAR 0.535 ± 0.193 0.143 ± 0.230 0.264 ± 0.283
Soft-Polarization-MNAR 0.546 ± 0.239 0.187 ± 0.245 0.432 ± 0.409
Latent-Factor-MNAR 0.358 ± 0.296 0.331 ± 0.312 0.301 ± 0.305
Cluster-MNAR 0.321 ± 0.277 0.317 ± 0.287 0.356 ± 0.311
Two-Phase-MNAR 0.521 ± 0.321 0.557 ± 0.313 0.218 ± 0.320
Overall 0.417 ± 0.314 0.378 ± 0.338 0.326 ± 0.343
Table 5: Imputation Accuracy ± Standard Deviation by Method (Zero-shot models)
Note that the normalized numbers are slightly different from Tab. 1 because we did not include every other zero-shot method in the normalization in Tab. 1.
Method MCAR NN-MNAR Self-Masking-MNAR Col-MAR
TabImpute+ 0.749 ± 0.246 0.847 ± 0.200 0.929 ± 0.178 0.883 ± 0.255
EWF-TabPFN 0.746 ± 0.247 0.841 ± 0.201 0.935 ± 0.179 0.882 ± 0.254
TabImpute 0.860 ± 0.222 0.828 ± 0.230 0.246 ± 0.254 0.551 ± 0.350
TabImpute (MCAR then MAR) 0.751 ± 0.284 0.666 ± 0.355 0.208 ± 0.243 0.499 ± 0.354
TabImpute (More Heads) 0.690 ± 0.261 0.700 ± 0.233 0.227 ± 0.255 0.523 ± 0.321
EWF-TabPFN Fine-Tuned 0.306 ± 0.189 0.279 ± 0.220 0.406 ± 0.313 0.339 ± 0.338
TabPFN 0.066 ± 0.193 0.067 ± 0.194 0.162 ± 0.224 0.160 ± 0.254
Method Block-MNAR Seq-MNAR Overall
TabImpute+ 0.804 ± 0.272 0.672 ± 0.361 0.814 ± 0.270
EWF-TabPFN 0.795 ± 0.279 0.670 ± 0.358 0.811 ± 0.271
TabImpute 0.778 ± 0.288 0.834 ± 0.214 0.683 ± 0.342
TabImpute (MCAR then MAR) 0.722 ± 0.287 0.805 ± 0.277 0.608 ± 0.362
TabImpute (More Heads) 0.668 ± 0.258 0.729 ± 0.260 0.589 ± 0.317
EWF-TabPFN Fine-Tuned 0.385 ± 0.305 0.347 ± 0.326 0.344 ± 0.288
TabPFN 0.048 ± 0.138 0.126 ± 0.262 0.105 ± 0.218
Table 6: Synthetic Data Generation Parameters
Missingness Pattern Parameter Name Symbol Value
MCAR Missing probability pp 0.4
Col-MAR Missing probability pp 0.4
NN-MNAR Neighborhood size $|N_{ij}|$ Variable
Network layers $L$ Random
Network depth $d$ Random
Weight initialization $W$ Random
Bias initialization $b$ Random
Self-Masking-MNAR Coefficient set $\alpha$ $\{-2,-1,1,2\}$
Target missing proportion $p_{missing}$ Variable
Block-MNAR Missing probability pp 0.4
Matrix size N NN 100
Matrix size T $T$ 50
Row blocks $B_{r}$ 10
Column blocks $B_{c}$ 10
Convolution type - mean
Seq-MNAR Missing probability pp 0.4
Algorithm - epsilon_greedy
Pooling - False
Epsilon $\epsilon$ 0.4
Epsilon decay $\gamma$ 0.99
Random seed $s$ 42
Panel-MNAR No explicit hyperparameters (dropout time sampled uniformly)
Polarization-MNAR Threshold quantile $q_{thresh}$ 0.25
Soft-Polarization-MNAR Polarization alpha $\alpha$ 2.5
Polarization epsilon $\epsilon$ 0.05
Latent-Factor-MNAR Latent rank (low) $k_{low}$ 1
Latent rank (high) $k_{high}$ 5
Cluster-MNAR Number of row clusters $K_{R}$ 5
Number of column clusters $K_{C}$ 4
Two-Phase-MNAR Cheap feature fraction $f_{cheap}$ 0.4
Censoring-MNAR Censoring quantile $q_{censor}$ 0.25
Table 7: OpenML datasets
Dataset Size Domain Description
EgyptianSkulls $150\times 5$ Anthropology Cranial measurements over time in Egypt
humans_numeric $75\times 15$ Biology Human body measurements
FacultySalaries $50\times 5$ Education/Economics University faculty salary data
SMSA $59\times 16$ Demographics/Economics U.S. metropolitan statistical area data
Student-Scores $56\times 13$ Education Student exam scores
analcatdata_election2000 $67\times 15$ Political science 2000 U.S. presidential election results
analcatdata_gviolence $74\times 9$ Criminology Gun violence statistics
analcatdata_olympic2000 $66\times 12$ Sports/Economics Olympic results and country stats
baskball $96\times 5$ Sports analytics Basketball performance data
visualizing_hamster $73\times 6$ Education/Toy Example dataset for teaching
witmer_census_1980 $50\times 5$ Demographics U.S. census microdata (1980)
MercuryinBass $53\times 10$ Environmental chemistry Mercury concentrations in fish
SolarPower $204\times 5$ Energy/Engineering Solar power output records
WineDataset $178\times 14$ Chemistry/Oenology Wine physicochemical properties
alcohol-qcm-sensor $125\times 15$ Analytical chemistry Alcohol detection sensor readings
benzo32 $195\times 33$ Chemistry/Toxicology Benzodiazepine compound data
machine_cpu $209\times 7$ Computer systems Predicting CPU performance
pwLinear $200\times 11$ Mathematics/Engineering Piecewise linear regression benchmark
pyrim $74\times 28$ Chemistry/Pharmacology Pyrimethamine bioassay compounds
slump $103\times 10$ Civil engineering Concrete slump test properties
ICU $200\times 20$ Medicine Intensive care patient data
appendicitis_test $106\times 8$ Medicine Appendicitis diagnosis
appendicitis_test_edsa $106\times 8$ Medicine Educational appendicitis dataset
breast-cancer-coimbra $116\times 10$ Medicine Breast cancer diagnosis data
Rainfall-in-Kerala-1901-2017 $117\times 18$ Climate science Rainfall time series in Kerala
pollution $60\times 16$ Environmental science Air pollution measurements
treepipit $86\times 10$ Ecology Bird habitat distribution
autoPrice $159\times 16$ Business/Economics Automobile pricing dataset
dataset_analcatdata_creditscore $100\times 7$ Finance Credit scoring dataset
Swiss-banknote-conterfeit-detection $200\times 7$ Finance/Fraud Banknote authenticity classification
Glass-Classification $214\times 10$ Forensics/Materials Glass chemical composition (forensics)
chatfield_4 $235\times 13$ Statistics/Time series Textbook time series data (Chatfield)
chscase_vine1 $52\times 10$ Agriculture/Statistics Vine growth study
edm $154\times 18$ Education Student learning performance
metafeatures $75\times 32$ Meta-learning Dataset-level features
rabe_131 $50\times 6$ Chemistry/Benchmark Spectroscopy regression dataset
rabe_148 $66\times 6$ Chemistry/Benchmark Spectroscopy regression dataset
rabe_265 $51\times 7$ Chemistry/Benchmark Spectroscopy regression dataset
sleuth_case1201 $50\times 7$ Statistics/Education Applied regression textbook data
sleuth_ex1605 $62\times 6$ Statistics/Education Applied regression textbook data
wisconsin $194\times 33$ Medicine Wisconsin breast cancer dataset
Table 8: Non-normalized RMSE values for MCAR pattern by dataset. Dataset columns are standardized based on observed values (mean 0, variance 1).
Dataset TabImpute+ HyperImpute MissForest OT
EgyptianSkulls 0.917 0.961 0.946 0.966
FacultySalaries 0.693 0.811 0.567 0.572
Glass-Classification 0.937 0.810 0.886 0.902
ICU 1.045 1.137 1.112 1.045
MercuryinBass 1.373 1.377 1.343 1.359
Rainfall-in-Kerala-1901-2017 0.942 0.902 0.950 0.958
SMSA 0.813 0.939 0.786 0.807
SolarPower 0.801 0.873 0.888 0.914
Student-Scores 0.391 0.441 0.435 0.477
Swiss-banknote-conterfeit-detection 0.976 0.856 0.730 0.786
WineDataset 0.928 0.803 0.797 0.789
alcohol-qcm-sensor 0.590 0.580 0.517 0.560
analcatdata_election2000 0.544 0.662 0.615 0.722
analcatdata_gviolence 0.795 0.801 0.877 0.794
analcatdata_olympic2000 1.867 1.922 1.895 1.945
appendicitis_test 0.860 0.650 0.792 0.753
appendicitis_test_edsa 0.668 0.532 0.569 0.586
autoPrice 0.834 0.762 0.702 0.734
baskball 1.027 1.100 1.109 1.017
benzo32 1.055 1.055 1.053 0.950
breast-cancer-coimbra 1.058 1.183 1.098 1.104
chatfield_4 0.475 0.531 0.492 0.522
chscase_vine1 0.808 0.996 0.833 0.945
dataset_analcatdata_creditscore 1.063 1.131 1.122 1.065
divorce_prediction 0.525 0.528 0.467 0.493
edm 0.763 0.553 0.566 0.570
humans_numeric 0.948 0.970 0.991 1.011
machine_cpu 1.037 0.956 0.858 0.962
metafeatures 2.263 2.090 2.225 2.263
pollution 1.967 2.065 1.837 1.945
pwLinear 1.075 1.356 1.356 1.092
pyrim 0.877 0.863 0.859 0.842
rabe_131 0.790 1.003 1.031 0.997
rabe_148 0.984 1.017 0.958 1.045
rabe_265 1.131 1.298 1.090 1.106
sleuth_case1201 1.054 0.939 0.963 0.998
sleuth_ex1605 1.087 1.257 1.071 0.969
slump 0.871 0.650 0.805 0.787
treepipit 1.264 1.169 1.172 1.167
visualizing_hamster 0.872 0.905 0.750 0.774
wisconsin 0.777 0.685 0.783 0.817
witmer_census_1980 0.704 0.616 0.637 0.653